OpenStudy (anonymous):

(10x3 - 8x2 - 12x) ÷ 2x) simplify using distributive property

OpenStudy (luigi0210):

\[\frac{ 10x^3-8x^2-12x }{ 2x }=\frac{ 2x(5x^2-4x-6) }{ 2x }\]

OpenStudy (anonymous):

so that's the answer??

OpenStudy (anonymous):

Im confused..

OpenStudy (anonymous):

the ans will be 5x^2 -4X-6......becos the 2X cancels

OpenStudy (anonymous):

okay thanks :)

OpenStudy (luigi0210):

-_-